OpenStudy (anonymous):

Check my answer? Question: Write an equation in slope-intercept form for the line parallel to y = 5x – 2 that passes through the point (8, –2). My answer: Okay so we know that our slope is 5, so now we can use the point slope formula. y - y1 = m (x - x1) = y - (- 2) = 5(x - 8) = y + 2 = 5x - 40, now we subtract 2 from both sides. y + 2 - 2 = 5x - 40 + 2 = y = 5x - 38.

hartnn (hartnn):

just 1 error.

hartnn (hartnn):

y+2 = 5x-40 subtract 2 from both sides. y+2-2 = 5x-40-2 -40-2 =... ?

OpenStudy (anonymous):

- 42

hartnn (hartnn):

so, y= 5x-42 is correct :)

OpenStudy (anonymous):

Thanks

hartnn (hartnn):

welcome ^_^